OpenStudy (anonymous):

A phone company charges $25 per month, plus an installation fee of $50. If you have paid the phone company $375, how many months have you had the phone?

OpenStudy (anonymous):

375 = 25*t + 50 where 't' is time in months solve for t

OpenStudy (anonymous):

...

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

Add 50 to both sides?

OpenStudy (anonymous):

nope, you want to subtract 50 from both sides 375 - 50 = 25*t + 50 - 50

OpenStudy (anonymous):

Ah. Okay. So you're left with 320=25t?

OpenStudy (anonymous):

not quite... 375 - 50 = 325.......... not 320 so you get 325 = 25t

OpenStudy (anonymous):

Lol. Yeah. So, Then, do you divide them both by 25?

OpenStudy (anonymous):

correct, and that should give you your answer

OpenStudy (anonymous):

t=13?

OpenStudy (anonymous):

yep, and t is the number of months

OpenStudy (anonymous):

thank you! (:

OpenStudy (anonymous):

@hadleydarcym you're welcome